# Conformalized Credal Regions for Classification with Ambiguous Ground Truth

### v1

@@ Appendix A Uncertainty Quantification
Other measures for AU and EU are also available in the context of credal regions [4, 27, 30], and they can be used in place of upper and lower entropy to quantify EU and AU within our credal region $\mathcal{P}$, as long as the measure chosen for the total uncertainty is bounded. We chose upper and lower entropy because of their ease of computation. In fact, let us give a case in which the quantities in (6) can be easily calculated or bounded. Let $\text{ex}\mathcal{P}$ denote the extreme elements of $\mathcal{P}$, that is, those elements that cannot be written as a convex combination of one another. We have that $\mathcal{P}=\text{Conv}(\text{ex}\mathcal{P})$, and $\text{Conv}(\cdot)$ denotes the convex hull operator. Then, the following was proven in [8, Theorem 8].
@@
Calculating $\underline{H}(P^{\text{ex}})$ and $\overline{H}(P^{\text{ex}})$ is immediate from (5). On the other hand, the supremum $\sup_{\beta\in\Delta^{S-1}}\sum_{s=1}^{S}\beta_{s}H(P^{\text{ex}}_{s})$ can be computed in polynomial time. [^8] We also point out how bounds for $\text{TU}(\mathcal{P})$, $\text{AU}(\mathcal{P})$, and $\text{EU}(\mathcal{P})$ in terms of the entropy of the lower probability and of the entropy of the upper probability, $H(\underline{P})$ and $H(\overline{P})$, respectively – not to be confused with the lower and upper entropy $\underline{H}(P)$ and $\overline{H}(P)$, respectively – were given in [6, Theorem 13].

### v2

@@ Appendix B Uncertainty Quantification
Other measures for AU and EU are also available in the context of credal regions (Bronevich & Rozenberg, 2021; Hofman et al., 2024; Hüllermeier & Waegeman, 2021), and they can be used in place of upper and lower entropy to quantify EU and AU within our credal region $\mathcal{P}$, as long as the measure chosen for the total uncertainty is bounded. We chose upper and lower entropy because of their ease of computation. In fact, let us give a case in which the quantities in equation 7 can be easily calculated or bounded. Let $\text{ex}\mathcal{P}$ denote the extreme elements of $\mathcal{P}$, that is, those elements that cannot be written as a convex combination of one another. We have that $\mathcal{P}=\text{Conv}(\text{ex}\mathcal{P})$, and $\text{Conv}(\cdot)$ denotes the convex hull operator. Then, the following was proven in (Caprio et al., 2024b, Theorem 8).
@@
Calculating $\underline{H}(P^{\text{ex}})$ and $\overline{H}(P^{\text{ex}})$ is immediate from equation 6. On the other hand, the supremum $\sup_{\beta\in\Delta^{S-1}}\sum_{s=1}^{S}\beta_{s}H(P^{\text{ex}}_{s})$ can be computed in polynomial time. [^8] We also point out how bounds for $\text{TU}(\mathcal{P})$, $\text{AU}(\mathcal{P})$, and $\text{EU}(\mathcal{P})$ in terms of the entropy of the lower probability and of the entropy of the upper probability, $H(\underline{P})$ and $H(\overline{P})$, respectively – not to be confused with the lower and upper entropy $\underline{H}(\mathcal{P})$ and $\overline{H}(\mathcal{P})$, respectively – were given in (Caprio et al., 2024a, Theorem 13).
Appendix C Dermatology Dataset Details
In Section 5, by studying the Dermatology DDX dataset, we follow (Liu et al., 2020; Stutz et al., 2023a; b; c) and consider a very ambiguous as well as safety-critical application in dermatology: skin condition classification from multiple images. Such a dataset is from Liu et al. (2020), and consists of $1949$ test examples and $419$ classes with up to $6$ color images resized to $448\times 448$ pixels. The classes, i.e., conditions, were annotated by various dermatologists who provide partial rankings. These rankings are aggregated deterministically to obtain the plausibilities $\lambda$ using the inverse rank normalization procedure of (Liu et al., 2020) described in (Stutz et al., 2023b, Section 3.1).
Appendix D Runtime of CIFAR-10H with More Classes
As a sanity check, for CIFAR-10H we computed the runtime for each testing point across different coverage levels using $5$ random seeds. As can be seen from the plot below, the average runtime for a $3$-class problem is about $1.5$ ms and $3.5$ for a $5$-class problem, which we deem to be a reasonable increase.